OpenStudy (anonymous):

Which expression is the simplified form of √k^21

OpenStudy (anonymous):

you know that (√k)^2=k right?

OpenStudy (anonymous):

\[\large (\sqrt{k})^{21}=(k^{\frac{1}{2}})^{21}=k^{10}k^{\frac{1}{2}}=k^{10}\sqrt{k} \]